### v4

# Three results on twisted $G$-codes and skew twisted $G$-codes

Abstract
In this paper, we solve an open question posed in the original paper on twisted skew group codes concerning the conditions under which a twisted skew group code is checkable. We also prove that all ideals of dimension three over a twisted group algebra are abelian group codes, thereby generalizing a previous result for group algebras. Finally, we establish a bound on the dimension and minimum distance of a twisted group code, and characterize the cases in which this bound is attained.
Key Words : Coding theory; group ring; twisted group ring; twisted skew group ring
1 Introduction
The foundational results in [12] and [10] started the research in Information Theory and Error Correcting code Theory. Since the early days of the field, several algebraic structures have been used to obtain linear codes with the necessary properties for each purpose.
Let $G$ be a group and $K$ be a commutative ring. Then the group algebra $KG$ is the formal linear combination of elements of $G$, where product is given by the group operation in $G$. This ring was initially studied as they have a strong connection with representation theory, where linear representation of a group are in bijection with modules over the group algebras.
In [2] Berman proved that binary Reed Muller codes are ideals in a binary group algebra over an abelian group. In the following decade some mathematicians studied ideals over group algebras as they provide a rich framework to obtain new linear codes with interesting properties.
This research area remains very active. For example, in [3], the authors present the first bound relating the minimum distance, dimension of a linear code, and the underlying group. Another result connecting ring theory and coding theory is [6], where it is shown that group codes of dimensions two and three are commutative.
Recently, new generalizations of group algebras have been used in coding theory, such as twisted group codes [7], skew group codes [5] or twisted skew G-codes [1]. These structures provide interesting examples of linear codes. To illustrate this, the ternary extended self-dual Golay code studied in [7] are twisted group codes, or in [1], the authors present a $6,3,4]_{9}$, where the authors point out that it is constructed in a simpler way that in [9]
In addition, there are still some open problems regarding the properties of such codes. In [1], the authors proved a sufficient condition for a twisted group code to be code checkable. However, they propose as a question if there is an equivalent of such result to the case of twisted skew group codes.
In this paper, some results regarding group codes or twisted group codes are generalized. To begin with, using an approach similar to that in the original paper, we give a positive answer to the question in [1] about checkable codes in twisted skew group rings. After this, we give a sufficient condition for a twisted code to be permutation equivalent to a group code, which allows us to prove that all twisted codes of dimension 2 and 3 are abelian, and that may be useful in future with new twisted codes problems. Finally, in [13] a bound is proved on the dimension and minimal distance of an ideal over a twisted group algebra, and in [4] group codes where the bound is reach are characterized. We present a generalization of this result to the case of twisted group codes.
2 Preliminary results
We start by recalling some basic background and the notation we will use throughout this work. For more information, see [11].
Definición 1.
A crossed system is a quartet $(G,B,\sigma,\alpha)$ where $G$ is a group, $B$ is an $R$-algebra, being $R$ a ring and $\sigma$, $\alpha$ are applications
$\sigma:G\rightarrow Aut(B)\hskip 28.45274pt\alpha:G\times G\rightarrow U(B)$ (1)
where $Aut(B)$ is the group of automorphisms of $B$ and $U(B)$ denotes the subgroup given by the units of $B$, and, if we write $\prescript{g}{}{{b}}=\sigma(g)(b)$, then the following identities must hold:
  1. 1.
    $\prescript{x}{}{{(\prescript{y}{}{{b}})}}=\alpha(x,y)\prescript{xy}{}{{b}}\alpha(x,y)^{-1}$$\forall b\in B,x,y\in G$
  2. 2.
    $\alpha(x,y)\alpha(xy,z)=\prescript{x}{}{\alpha(y,z)}\alpha(x,yz)$ $\forall x,y,z\in G$
  3. 3.
    $\alpha(x,e)=\alpha(e,x)=1$ $\forall x\in G$
A map $\alpha$ that holds the previous conditions is called a 2cocycle2-cocycle.
For every crossed system we can define what is known as a crossed product or twisted skew group ring.
Definición 2.
Let $(G,B,\sigma,\alpha)$ be a crossed system. The crossed product, also called twisted skew group ring, of $B$ over $G$, $B^{\alpha}[G;\sigma]$, is the free $B$-module generated by the elements $g\in G$, denoted by $\overline{g}$, with the inner operation
$(a\overline{x})\cdot(b\overline{y})=a\prescript{x}{}{b}\alpha(x,y)\overline{xy}\forall a,b\in B,x,y\in G$ (2)
and which is extended by linearity.
In this work, we will deal with a twisted skew group ring over a finite field.
3 Code Checkable
Let $G$ be a group, let $p$ be a prime number, and $\mathbb{F}_{p^{n}}$ be the finite field of $p^{n}$ elements, and let $(G,\mathbb{F}_{p^{n}},\sigma,\alpha)$ be a crossed system. Then, we know that $Aut(\mathbb{F}_{p^{n}})=Gal(\mathbb{F}_{p},\mathbb{F}_{p^{n}})\simeq C_{n}$ the cyclic group of order $n$, generated by the Frobenious map. In particular, we have that $\mathbb{F}_{p}$ is fixed by the action of $G$ and $\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ is an algebra over its prime field.
Let $\hat{G}=\mathbb{F}_{p^{n}}^{*}\times G$ and let
$\hat{G}\times\hat{G}$ $\longrightarrow\hat{G}$
$((a,g),(b,h))$ $\longmapsto(a\prescript{g}{}{b}\alpha(g,h),gh)$
Proposition 1.
$\hat{G}$ with the previous operation is a group.
Proof.
It is close under the operation as $\mathbb{F}_{p^{n}}$ is an integral domain, and the associativity comes from the associativity of the product in a twisted skew group ring.
It has an identity element, given by $1_{\hat{G}}=(1_{\mathbb{F}_{p}^{n}},1_{G})$ as for all $(x,y)\in\hat{G}$
$(x,y)1_{\hat{G}}$= $(x,y)(1_{\mathbb{F}_{p}^{n}},1_{G})$
$=$ $(x\prescript{y}{}{1_{\mathbb{F}_{p}^{n}}},y1_{G})$
$=$ $(x,y)$
$=$ $(1_{\mathbb{F}_{p}^{n}}\prescript{1_{G}}{}{x},1_{G}y)$
$=$ $(1_{\mathbb{F}_{p}^{n}},1_{G})(x,y)$
$=$ $1_{\hat{G}}(x,y)$
Finally, the inverse element of $(x,y)\in\hat{G}$ is $(\prescript{y^{-1}}{}{}\left(\frac{1}{x\alpha(y,y^{-1})}\right),y^{-1})$, as
$(x,y)(\prescript{y^{-1}}{}{}\left(\frac{1}{x\alpha(y,y^{-1})}\right),y^{-1})=\left(\frac{x\alpha(y,y^{-1})}{x\alpha(y,y^{-1})},yy^{-1}\right)=(1_{\mathbb{F}_{p^{n}}},1_{G})=1_{\hat{G}}$ (3)
and the other side is proven analogously. ∎
Proposition 2.
There is an epimorphism of $\mathbb{F}_{p}$-algebras over $\mathbb{F}_{p}\hat{G}\longrightarrow\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$.
Proof.
The epimorphism $\psi:\mathbb{F}_{p}\hat{G}\longrightarrow\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ given by $\psi(a\overline{(b,c)})=ab\overline{c}$ for all $a\in\mathbb{F}_{p}$, $(b,c)\in\hat{G}$ and extended by linearity. It respects addition by definition, so we only have to prove that it respects product, and to do so we can focus on the basis elements
$\psi(\overline{(a,g)}\overline{(b,h)})=\psi((a\prescript{g}{}{b}\alpha(g,h),gh)=a\prescript{g}{}{b}\alpha(g,h)\overline{gh}=(a\overline{g})(b\overline{h})=\psi(\overline{(a,g)})\psi(\overline{(b,h)}$ (4)
To begin with, we recall the definition of checkable ideals.
Definición 3.
A right ideal $I\leq A$ is called checkable if there exists an element $v\in A$ such that
$I=\{\,a\in A\mid va=0\,\}=\operatorname{Ann}_{r}(v)=\operatorname{Ann}_{r}(Av).$
Here, checkable left ideals are defined analogously via the left annihilator of a principal right ideal. A group algebra $K^{\alpha}G$ is called code-checkable if all right ideals of $K^{\alpha}G$ are checkable.
In [3] the following result is proved.
Corollary 3 (Corollary 3.2, [3]).
Let $\operatorname{char}K=p$ and let $B_{0}(G)$ be the principal $p$-block of $KG$. The following are equivalent:
  1. 1.
    $G$ is $p$-nilpotent with a cyclic Sylow $p$-subgroup.
  2. 2.
    KGKG is a code-checkable group algebra.
  3. 3.
    All right ideals in $B_{0}(G)$ are checkable.
In [1] a question regarding when twisted-skew G-code are code checkable is proposed.
Proposition 4 ([1, Proposition 1.10]).
Let $\operatorname{char}K=p$. If $G$ is $p$-nilpotent with a cyclic Sylow $p$-subgroup, then every left (right) ideal in a twisted group algebra is principal.
Question 1 ([1, Question 1.11]).
Is there an extension of Proposition 1.10 to left ideals in twisted skew group rings?
To present the answer, we begin with some previous results. First of all, note that for all $(a,g)\in\hat{G}$ we have that
$(a,g)^{b}=(\prod_{k=0}^{b-1}\alpha(g,g^{k})\prescript{g^{k}}{}{a},g^{b})$ (5)
Lemma 5.
$G$ is $p$-nilpotent if and only if $\hat{G}$ is $p$-nilpotent
Proof.
$(\Longrightarrow)$
Suppose that $G$ is $p$-nilpotent, then all elements of order coprime with $p$ form a subgroup $H\leq G$. Then the set $\hat{H}=\mathbb{F}_{p^{n}}^{*}\times H$ is precisely the set of all elements whose order is coprime $p$ of $\hat{G}$.
First, let $(a,b)\in\hat{G}$ of order a multiple of $p$ i.e. $(a,b)=p^{c}h$ with $gcd(h,p)=1$. Then $(a,b)^{h}=(a^{\prime},b^{h})$ is of order $p^{c}$, so $b^{h}$ is of order a power of $p$, and $b\not\in H$.
In addition, if $(a,b)\in\hat{G}$ is of order coprime with $p$, by the same argument $b$ is of order coprime with $p$ and $(a,b)\in\hat{H}$.
We have shown that $\hat{H}=\mathbb{F}_{p^{n}}^{*}\times H$ is the set of all elements of order coprime with $p$ of $\hat{G}$. The fact that it is a subgroup is proven analogously to 1 using the fact that $H$ is a subgroup.
[…]
Let $\pi:\hat{G}\longrightarrow G$ the projection on the second component. It is clearly a epimorphism of groups. Let $\hat{H}$ be the group form by all elements of order coprime with $p$. We will show that $\pi(\hat{H})=H\leq G$ is a group form by all elements whose order coprime with $p$ in $G$.
First of all, as $\hat{H}$ is a subgroup, $H$ is also a group, so we only have to prove the $p$-nilpotent part. If $a\in G$ has order corpime with $p$, then $(1,a)$ has order coprime with $p$, so $a=\pi((1,a))\in\pi(\hat{H})=H$. Now, if $a\in H$, then it is the image of $(1,a)\in\hat{H}$, and $ord(a)=ord((1,a))$ so it has order coprime with $p$. ∎
Lemma 6.
$G$ has a cyclic Sylow $p$-subgroup if and only if $\hat{G}$ also.
Proof.
[…]
$(\Longrightarrow)$
Suppose that $G$ has a cyclic Sylow $p$-subgroup $P$. Then all $p$-subgroups of $G$ are cyclic from the first Sylow theorem. Now, let $\hat{P}$ be the $p$-Sylow subgroup of $\hat{G}$.
Let $(a,g),(b,g)\in\hat{P}$ with $a\not=b\in\mathbb{F}_{p^{n}}$. Then $(a,g)(b,g)^{-1}\in\hat{P}$, so
$(a,g)(b,g)^{-1}=(a,g)\left(\prescript{g^{-1}}{}{\left(\frac{1}{b\alpha(g,g^{-1})}\right)},g^{-1}\right)=(ab^{-1},1_{G})$
So $(ab^{-1},1_{G})\in\hat{P}$, but $ord((ab^{-1},1_{G}))$ is coprime with $p$, so the only possibility is $ord((ab^{-1},1_{G}))=1$ so $a=b$
[…]
Proposition 7.
If $G$ is $p$-nilpotent with a cyclic Sylow $p$-subgroup then The twisted group ring $\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ is code checkable
Proof.
$(\Longrightarrow)$
If $G$ is $p$-nilpotent with a cyclic Sylow $p$-subgroup then $\hat{G}$ is $p$-nilpotent with a cyclic Sylow $p$-subgroup. Then $\mathbb{F}_{p}\hat{G}$ is code checkable. Then let $I\trianglelefteq_{r}\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ a right ideal, and let $\hat{I}=\pi^{-1}(I)$. Then $\hat{I}\trianglelefteq_{r}\mathbb{F}_{p}\hat{G}$ and therefore is code checkable, so $\hat{I}=\operatorname{Ann}_{r}(v)$ with $v\in\mathbb{F}_{p}\hat{G}$. Therefore $I=\pi(\hat{I})=\pi(\operatorname{Ann}_{r}(v))=\operatorname{Ann}_{r}(\pi(v))$
This, in addition, also gives sufficient conditions for $\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ to have only principals left ideals. To do so, recall the well known result about Frobenius algebras.
Lemma 8.
Let $A$ be a Frobenius algebra and let $L\leq A$ be a left ideal. Set $I=\operatorname{Ann}_{r}(L)$. Then
$\operatorname{Ann}_{\ell}(I)=L.$
Now we can prove the generalization to our case.
Proposition 9.
Let $K^{\alpha}G$ be a code-checkable group algebra. Then every left ideal of $K^{\alpha}G$ is principal.
Proof.
Since $K^{\alpha}G$ is code-checkable, every right ideal is checkable by definition. By the previous result, in a Frobenius algebra, a right ideal $I$ is checkable if and only if its left annihilator $\operatorname{Ann}_{\ell}(I)$ is a principal left ideal. Hence
$\operatorname{Ann}_{\ell}(I)\ \text{is principal for every right ideal }I\leq KG.$
Now let $L\leq KG$ be any left ideal. Consider the right ideal
$I:=\operatorname{Ann}_{r}(L).$
By the double annihilator property for Frobenius algebras (Proposition 2.2), we have
$L=\operatorname{Ann}_{\ell}(\operatorname{Ann}_{r}(L))=\operatorname{Ann}_{\ell}(I).$
Thus $L$ is the left annihilator of a right ideal. By $(\ast)$, $\operatorname{Ann}_{\ell}(I)$ is principal, and therefore $L$ is principal.
Hence every left ideal of $K^{\alpha}G$ is principal as in [1] it is proven that it is Frobenius. ∎
4 Abelian ideals in twisted group algebra
First of all, recall some basic definitions of coding theory.
Definición 4.
Let $C_{1},C_{2}\subset\mathbb{F}_{q}^{m}$ codes are said to be equivalent if there exist a diagonal matrix $D=diag(\lambda_{1},\cdots,\lambda_{m})\in Mat_{m}(\mathbb{F}_{q})$ and a permutation matrix $P$ such that $C_{1}=\{PDx;x\in C_{2}\}$. They are said to be permutation equivalent if $D=Id_{m}$.
In [6] the following result regarding when ideals over group algebra are permutation equivalent to abelian group algebras.
Teorema 10 (Theorem 1, [6]).
Let $C$ be a $G$-code over a finite field $K$ for a finite group $G$. If
$\dim_{K}(C)\leq 3,$
then $C$ is permutation equivalent to an abelian group code.
We will generalize the result to the case of twisted group codes. To begin with, we need some previous results, for which in the same paper appears
Lemma 11 (Lemma 1.3, [6]).
Let $K$ be a field and let $G$ and $H$ be two groups of the same order $n<\infty$. Suppose that there exist two normal subgroups $N\triangleleft G$ and $F\triangleleft H$ such that $G/N\cong H/F$. If $N$ acts trivially on some (left, right, or two-sided) ideal $I\subseteq KG$, then $I$ is permutation equivalent to some (left, right, or two-sided) ideal of the ring $KH$.
Our generalization needs the following definition
Definición 5.
Let $K^{\alpha}G$ be a twisted group algebra with $G$ finite. We will say that a subgroup $N$ of $G$ acts by scalar on a subset $S\subset K^{\alpha}G$ if there exists a $\alpha$- homomorphism $\lambda:N\longrightarrow K$ such that
$\overline{u}x=\lambda(u)x\quad\forall x\in S,u\in N$ (5)
Now we can state that
Lemma 12.
Let $K^{\alpha}G$ be a twisted group algebra with $G$ finite. Let $H$ be a group of the same order as $G$. Suppose that there exist two normal subgroups $N\triangleleft G$ and $F\triangleleft H$ such that $G/N\cong H/F$. If $N$ acts by scalar on some (left, right, or two-sided) ideal $I\subseteq K^{\alpha}G$, then $I$ is equivalent to some (left, right, or two-sided) ideal of the ring $KH$.
Proof.
Let $s=|G/N|$ and denote by $g_{1},\ldots,g_{s}$ a complete set of representatives of $G$ modulo $N$. Thus
$G/N=\{g_{1}N,\ldots,g_{s}N\}\quad\text{and}\quad G=\bigcup_{i=1}^{s}g_{i}N.$
Fixing the isomorphism $f\colon G/N\to H/F$, we can choose a representative system $\{h_{i}\}_{i=1}^{s}$ of the group $H$ modulo $F$ such that
$f(g_{i}N)=h_{i}K,\quad i=1,\ldots,s.$
For each pair of indices $i,j\in\{1,\ldots,s\}$, let $k(i,j)$ be defined by the equality
$g_{i}Ng_{j}N=g_{k(i,j)}N.$
Then we have
$g_{i}g_{j}=g_{k(i,j)}u_{ij}\quad\text{for some }u_{ij}\in N,$
and similarly
$h_{i}h_{j}=h_{k(i,j)}v_{ij}\quad\text{for some }v_{ij}\in F.$
Since $|N|=|F|=n/s$, we can fix a one-to-one mapping
$\tau\colon N\to F.$
Define $\varphi\colon G\to H$ as follows: for an arbitrary $x\in G$, the element $x$ belongs to exactly one coset $Ng_{i}$ of $G$ modulo $N$. Set
$\varphi(x)=\frac{\alpha(xg_{i}^{-1},g_{i}x^{-1})}{\alpha(xg_{i}^{-1},g_{i})\lambda(xg_{i}^{-1})}\overline{h_{i}\,\tau(xg_{i}^{-1})}$
Clearly, $\varphi$ is a well-defined one-to-one map.
Let $x=\sum_{g\in G}a_{g}\overline{g}\in I$ and let $\lambda$ be associated $\alpha$-homomorphism of $N$ over $I$. Then, for all $u\in N$ we have that
$\lambda(u)\sum_{g\in G}a_{g}\overline{g}=\lambda(u)x=\overline{u}x=\sum_{g\in G}a_{g}\alpha(u,g)\overline{ug}$ (6)
So $a_{g}=\frac{1}{\lambda(u)}a_{u^{-1}g}\alpha(u,u^{-1}g)$, or equivalently $a_{g}=\frac{1}{\lambda(u^{-1})}a_{ug}\alpha(u^{-1},ug)$. In addition, we have that
$\alpha(u^{-1},ug)=\frac{\alpha(u^{-1},u)}{\alpha(u,g)}$ (7)
and therefore we have the equality
$a_{ug}=\lambda(u)\frac{\alpha(u,g)}{\alpha(u^{-1},u)}a_{g}$ (8)
Then $x=\sum_{i=1}^{s}a_{g_{i}}\sum_{u\in N}\frac{\lambda(u)\alpha(u,g)}{\alpha(u^{-1},u)}\overline{ug_{i}}$. Now, note that for all $i$ we have that
$\varphi\left(\sum_{u\in N}\frac{\lambda(u)\alpha(u,g)}{\alpha(u^{-1},u)}\overline{ug_{i}}\right)$ $=\sum_{u\in N}\frac{\lambda(u)\alpha(u,g)}{\alpha(u^{-1},u)}\varphi(\overline{ug_{i}})$
$=\sum_{u\in N}\frac{\lambda(u)\alpha(u,g)}{\alpha(u^{-1},u)}\frac{\alpha(u^{-1},u)}{\alpha(u,g)\lambda(u)}\overline{h_{i}\tau(u)}$
$=h_{i}\sum_{u\in N}\tau(u)$
$=h_{k}F_{\sum}$
Also, for all $h\in H$, we have that $h=h_{j}v$ for a unique $h_{j}$ and $v\in F$. We have that
$\overline{h}\varphi(x)$ $=\overline{h_{j}v}\sum_{i=1}^{s}a_{g_{i}}\overline{h_{i}}F_{\sum}$
$=\sum_{i=1}^{s}a_{g_{i}}\overline{h_{j}v}\overline{h_{i}}F_{\sum}$
$=\sum_{i=1}^{s}a_{g_{i}}\overline{h_{j}v}\overline{h_{i}}F_{\sum}$
$=\sum_{i=1}^{s}a_{g_{i}}\overline{h_{j}h_{j}v^{\prime}}F_{\sum}$
$=\sum_{i=1}^{s}a_{g_{i}}\overline{h_{k(j,i)}v^{\prime\prime}v^{\prime}}F_{\sum}$
$=\sum_{i=1}^{s}a_{g_{i}}\overline{h_{k(j,i)}}F_{\sum}$
$=\sum_{i=1}^{s}a_{g_{i}}F_{\sum}\overline{h_{k(j,i)}}$
$=\varphi\left(\sum_{i=1}^{s}a_{g_{i}}\sum_{u\in N}\frac{\alpha(u,g_{k(j,i)})}{\alpha(u^{-1},u)\lambda(u)}\overline{ug_{k(j,i)}}\right)\in\varphi(I)$
a similar calculation lead to $\varphi(x)\overline{h}\in\varphi(I)$ and we have the result
Now, we will start the proof of the following theorem. r
Teorema 13.
Let $K^{\alpha}G$ be a twisted group algebra whit $G$ finite. Let $C$ be a twisted $G$-code over a finite field $K$ for a finite group $G$. If
$\dim_{K}(C)\leq 3,$
then $C$ is an abelian group code.
We first start with dimension $1$.
Teorema 14.
Let $K^{\alpha}G$ be a twisted group algebra whit $G$ finite. If there exists a $1$ dimensional ideal then $K^{\alpha}G\simeq KG$. In particular, all $1$ dimensional ideal are permutation equivalent to an abelian group code.
Proof.
If there exist $I\triangleleft K^{\alpha}G$, then $I=Kv$ with $v\in K^{\alpha}G$. Then, for all $g\in G$, we have that $\overline{g}v=\lambda(g)v$ with $\lambda(g)\in K^{*}$. Therefore, we have
$\lambda(gh)\alpha(g,h)v=\alpha(g,h)\overline{gh}v=(\overline{g}\cdot\overline{h})=\overline{g}(\overline{h}v)=\lambda(g)\lambda(h)v$ (9)
And therefore
$\alpha(g,h)=\lambda(g)\lambda(h)\lambda(gh)^{-1}$ (10)
so $\alpha$ is a 22-coboundary and therefore $K^{\alpha}G\simeq KG$
This will help us to prove that
Teorema 15.
Let $K^{\alpha}G$ be a twisted group algebra whit $G$ finite, and let $I\triangleleft K^{\alpha}G$ of dimension $2$. Then $I$ is equivalent to some ideal in a commutative group ring over $K$ .
Proof.
If $I$ is not simple, then there exist a one dimensional ideal over $K^{\alpha}G$, and due to the previous lemma we have that $K^{\alpha}G\simeq KG$ and the result comes form the non twisted case.
Then by Schur’s lemma its endomorphism ring is a division ring $D$, that, as $K$ and $G$ is finite, $I$ is finite and $D$ is a field. Therefore, multiplication by $K^{\alpha}G$ defines an $K$-homomorphism over $D$, which we will denote by $\psi$.
As $D$ is commutative, $G^{\prime}$ acts by scalar. To see this, note that $x=\overline{g}\cdot\overline{h}\cdot\overline{g^{-1}}\cdot\overline{h^{-1}}$ satisfy $\psi(x)=\psi(g)\psi(h)\psi(g^{-1})\psi(h^{-1})=\psi(g)\psi(h)\psi(g)^{-1}\psi(h)^{-1}=Id$, and
$\psi(x)=\psi(\alpha(g,h)\alpha(gh,g^{-1})\alpha(ghg,h^{-1})\psi(\overline{ghg^{-1}h^{-1})}$
so we have that
$\psi\left(\overline{ghg^{-1}h^{-1}}\right)(a)=\frac{1}{\alpha(g,h)\alpha(gh,g^{-1})\alpha(ghg,h^{-1})}a$
and $G^{\prime}$ acts by scalar over $I$. Now, if we take $H=G^{\prime}\times G/G^{\prime}$, $N=F=G^{\prime}$ we are under the conditions for lemma 12
Now, for dimension $3$ we need the following lemma from [6]
Lemma 16 (Lemma 2.3 [6]).
Let $R$ be an $F$-algebra. Suppose that $I\triangleleft R$ and $\dim_{F}(I)=3$. If $M$ is a two-dimensional simple submodule of $IR$, then $M$ is a fully characteristic submodule of $IR$ (i.e. $f(M)\subseteq M$ for any $f\in\operatorname{End}(IR)$), in particular, $M\subseteq R$.
If $IR/N$ is a two-dimensional simple factor module of $IR$ for some submodule $N$, then $N$ is a fully characteristic submodule of $IR$, in particular, $N\triangleleft R$.
Now, we can prove
Teorema 17.
Let $K^{\alpha}G$ be a twisted group ring with $G$ finite and let $I\triangleleft K^{\alpha}G$ be an ideal of dimension $3$ over $K$. Then $I$ is equivalent to an abelian group code.
Proof.
If $I$ is simple, then the same argument as in dim $2$ remain valid. If $I$ has a 1 dimensional ideal then it is also true as then $K^{\alpha}G\simeq KG$ and the group ring case prove it. So it remain to prove the case when it has a $2$ dimensional sub ideal.
If $I$ has a 22-dimensional sub ideal $M\triangleleft I$, let $v_{1},v_{2},v_{3}\in I$ be a base of $I$ with $v_{1},v_{2}\in M$. Then we have that $\overline{g}v_{1},\overline{g}v_{2}\in M$ for all $g\in G$ and $\overline{g}v_{3}=\beta_{1}(g)v_{1}+\beta_{2}(g)v_{2}+\beta_{3}(g)v_{3}$. If we see the component of $v_{3}$ we have that $\alpha(g,h)\overline{gh}v_{3}=\overline{g}\cdot\overline{h}v_{3}$ so $\alpha(g,h)\beta_{3}(gh)=\beta_{3}(g)\beta_{3}(h)$ with $\beta_{3}\in\mathbb{F}_{p}^{n}$ for all $g,h\in G$.
If $\beta_{3}\not=0$ for all $g$ then $\alpha$ is a 22-coboundary and $K^{\alpha}G\simeq KG$ so the statement is true. Now, if $\beta_{3}(g)=0$ for some $g$, then as $\alpha(g,h)\beta_{3}(gh)=\beta_{3}(g)\beta_{3}(h)=0$ and $\alpha(g,h)\not=0$ we have that $\beta_{3}(gh)=0$ for all $h$, so $\beta_{3}(h)=0$ for all $h\in G$, so $1v_{3}\in<v_{1},v_{2}>=M$ absurd as they are linearly independent ∎
5 Bounds on distance and dimension of a group code
The next result was first proved for group codes in [4]. An extension to twisted group codes has been given proved in [13], and for twisted skew group ring, and the presented version in this paper, in [1].
Teorema 18 (Theorem 2.4, [1]).
If $0\neq C\leq K^{\alpha}[G,\sigma]=R$ is a twisted skew group code of minimum distance $d(C)$, then
$|G|\leq d(C)\cdot\dim C.$
In addition, in [4] group codes where the equality hold have been charactericed.
Teorema 19 (Theorem 2.10., [4]).
A $G$-code $C$ satisfies $d(C)\cdot\dim C=|G|$ if and only if there exist $H\leq G$ and $c\in KH$ such that $|H|=d(C)$, cKHcKH has dimension $1$, and $C=cKG$.
Now, we will generalize their result to the case of twisted group algebras. Following [8] we have
Definición 6.
Let $G$ be a group, and $S$ be a nonempty subset of $G$. A sequence $g_{1},\dots,g_{t}$ in $G$ has right $S$-rank $t$ if
$Sg_{i}:=\{sg_{i}\mid s\in S\}$
is not contained in $\bigcup_{j<i}Sg_{j}$ for all $i\in\{2,\dots,t\}$.
Let $R$ be a finite $G$-graded $K$-algebra with $G$ a finite group. For any $f\in R$, we denote by $T_{f}\colon C\to C$ the map $v\mapsto fv$. Then the following results from [4] are true in our case
Lemma 20.
[Lemma 2.2, [4]] Let $0\neq f\in R$ and $S=\operatorname{supp}(f)$. If there exists a sequence in $G$ with right $S$-rank $t$, then
$\dim fR=\operatorname{rank}_{K}(T_{f})\geq t.$
Teorema 21.
[Lemma 2.4, [4]] For any $f\in R$, define $T_{f}\colon R\to R$ by $v\mapsto fv$. Then
$|\operatorname{supp}(f)|\cdot\operatorname{rank}_{K}(T_{f})\geq|G|.$
Corollary 22.
For any nonzero ideal $C\triangleleft R$, we have
$d(C)\cdot\dim C\geq|G|.$
In particular,
$2\sqrt{|G|}\leq d(C)+\dim C\leq|G|+1.$
Proof.
Let $f\in C$ be such that $d(C)=|\operatorname{supp}(f)|$. Then the subcode $fR$ has the same minimum distance as $C$, but possibly a smaller dimension. Since $\dim(fR)=\operatorname{rank}_{K}(T_{f})$, the first claim follows from Theorem 22. The second claim is a consequence of the AM–GM inequality and the Singleton bound. ∎
Note that group ring, twisted group ring and twisted skew group ring are $G$-graded $K$-algebras, so all the previous results also apply to our case. In particular
Lemma 23.
If $R=\otimes_{g\in G}R_{g}$ is a strongly $G$-graded $K$-algebra with $\dim_{K}R_{g}=1$, then $R\simeq K^{\alpha}G$ for some 22-cocycle $\alpha:G\times G\longrightarrow K$.
So the interesting case for our results in a code perspective are all twisted group rings. Now we can prove the following result
Teorema 24.
Let $C$ be a twisted $G$-code
$d(C)\cdot\dim C=|G|$
if and only if there exist $H\leq G$ and $c\in R|_{H}$ such that $|H|=d(C)$, $cK^{\alpha}H$ has dimension $1$, and $C=cK^{\alpha}G$.
Proof.
Assume that there exist a subgroup $H\leq G$ and an element $c\in K^{\alpha}H$ satisfying the hypotheses of the statement. Then $C$ is the module induced from $cK^{\alpha}H$, regarded as a $K^{\alpha}H$-module, and it splits as a direct sum of precisely $[G:H]$ direct summands. Note that $\dim_{K}K^{\alpha}H=\dim_{K}K^{\alpha}Hg$ for all $g\in G$ so each summand is of dimension $1$.
Thus $\dim C=[G:H]$ and
$d(C)\cdot\dim C=|H|\cdot[G:H]=|G|.$
To prove the other direction, suppose that $d(C)\cdot\dim C=|G|$. Let $c\in C$ be an element such that $\operatorname{wt}(c)=d(C)$ and consider
$C_{0}:=cK^{\alpha}G\leq C.$
We have that $d(C_{0})=d(C)$, and by Corollary 22
$|G|\leq d(C_{0})\cdot\dim C_{0}=d(C)\cdot\dim C_{0}\leq d(C)\cdot\dim C=|G|.$
It follows that equality holds, and in particular $\dim C_{0}=\dim C$. Therefore
$C=C_{0}=cK^{\alpha}G.$
Now we will prove that there exist such $H$ as the theorem state. Let $H:=\operatorname{supp}(c)$, so that $|H|=d(C)$. Following the arguments and notation from the proof of Theorem 22 with $f=c$, we obtain
$t=\frac{|G|}{|H|}\quad\text{and}\quad G=Hg_{1}\,\dot{\cup}\,\cdots\,\dot{\cup}\,Hg_{t}$
for some $g_{1},\dots,g_{t}\in G$.
Replacing $c$ by the minimum weight codeword $ch^{-1}\in C$ for some $h\in\operatorname{supp}(c)$, we may assume that $1\in H$. Then for any $h\in H$ we have $h\in H\cap Hh$. We claim that this implies $H=Hh$.
Suppose by contradiction that $Hh\nsubseteq H$. Choose a maximal sequence $h_{1},\dots,h_{r}\in G$ with $r\geq 2$ such that $h_{1}=1$, $h_{2}=h$, and
$Hh_{i}\nsubseteq\bigcup_{1\leq j<i}Hh_{j}\quad\text{for }2\leq i\leq r.$
By maximality,
$G=\bigcup_{1\leq j\leq r}Hh_{j}.$
By definition of $t$ as in the proof of Theorem 22, we have $r\leq t$. However, since $H\cap Hh\neq\varnothing$, we obtain
$r>\frac{|G|}{|H|}=t,$
a contradiction, so $Hh=H$ for all $h\in H$, and therefore $H\cdot H\subseteq H$. Since $H$ is finite, this implies that $H$ is a subgroup of $G$ and $t=[G:H]$.
Finally, using the decomposition
$C=cK^{\alpha}G=cK^{\alpha}\Bigl(\bigsqcup_{i=1}^{t}Hg_{i}\Bigr)=\bigoplus_{i=1}^{t}(cK^{\alpha}Hg_{i}),$
in addition, as $\dim_{K}(cK^{\alpha}H)=\dim_{K}(cK^{\alpha}Hg_{i})$ because $\overline{g_{i}}$ is a unit, we obtain
$\dim C=t\cdot\dim(cK^{\alpha}H).$
By assumption,
$\dim C=\frac{|G|}{d(C)}=t,$
and hence $\dim(cK^{\alpha}H)=1$. This proves the claim. ∎
6 Conclusion
We have solved an open question concerning when a twisted skew $G$-code is code checkable. We have also generalized several results about group ring to the case of twisted group ring, such as that all twisted $G$-codes of dimension up to 3 are equivalent to an abelian group code, and also generalize a bound on the dimension and the minimum distance of a twisted $G$-code.
7 Acknowledgments
This research was supported by the Spanish Ministry of Science, Innovation and Universities under the FPU 2023 grant program.